OpenStudy (anonymous):

What is the sum of the geometric sequence –1, 6, –36, … if there are 7 terms?

OpenStudy (anonymous):

here, common ratio (r)= 6/-1=-6 and first term (a1)=-1 no of terms(n)=7 so, sum = a1(r^n-1)/(r-1) apply this and find the answer

OpenStudy (anonymous):

\[S_{n} = \frac{a_{1}(1 - r^{n})}{1 - r}\]

OpenStudy (anonymous):

Where n = 7 and a_1 = -1

OpenStudy (campbell_st):

the common ratio r = -6 n = 7 and a = -1 \[S _{7} = \frac{-1( (-6)^7 -1)}{-6 -1}\]

OpenStudy (anonymous):

sum =-39991

OpenStudy (anonymous):

thanks!!

OpenStudy (campbell_st):

yes thats it